XYZ stock has a current market price of $47.60 a share. the one-year call on XYZ stock with a strike price of $45 is priced at $3.20 whole the one-year put with a strike price of $45 is priced at $0.15. What is the risk-free rate of return?
Answer #1

As per Put Call Parity,

C + X/(1 + R)t = P + S

3.20 + 45/(1 + r) = 0.15 + 47.60

r = 1.01%

So,

Risk Free Rate = 1.01%